# DL-Chain: Scalable and Stable Blockchain Sharding with High Concurrency via Dual-Layer Consensus

### v1

@@ 7.1 Implementation
We developed a DL-Chain prototype in Golang, consisting of 1.4K lines of codes based on Harmony [6], a prominent permissionless blockchain sharding project. The FBFT consensus protocol, proposed by Harmony, is used as our intra-shard consensus to ensure a fair comparison to demonstrate DL-Chain’s performance improvements. Harmony, a traditional blockchain sharding unable to tolerate corrupted shards, serves as our baseline protocol. We also implemented a GearBox (CCS 22) [17] prototype as a state-of-the-art comparison. GearBox initially sets shards to a small size, prone to losing liveness. Upon detecting a corrupted shard, it reorganizes it until liveness is restored.
@@
7.2 Experimental Setup
We use a large-scale network of 2,550 nodes on 12 Amazon EC2 instances for testing. Each instance has a 128-core CPU and 50 Gbps network bandwidth, hosting up to 213 nodes. Docker containers and Linux Traffic Control managed inter-node communication, enforcing a 100 ms message delay and a 50 Mbps bandwidth limit per node. The experiment uses 512-byte transactions, and each block accommodates up to 4,096 transactions (i.e., 2MB blocks) as in existing work [36]. The transactions are based on historical data of Ethereum [38], in which the proportion of cross-shard transactions increases with the number of shards. The total fraction of malicious nodes is set at $f=1/4$, typical in practical network environments [25].
7.3 Parameter Settings
In DL-Chain, proposer shard and finalizer committee sizes should be adjusted to guarantee a negligible system failure probability, as mentioned in Section 6.1. In the baseline (Harmony), we leverage the equations based on classical hypergeometric distribution in [36] to determine shard sizes. We leverage the function proposed by GearBox and conduct 10 million simulations from the beginning of the epoch to get the average number of shards for implementing GearBox. The negligible failure probability is less than $2^{-17}\approx 7.6\times 10^{-6}$, the same as many existing works [19, 23], meaning that one failure occurs in about 359 years.
The parameter settings are shown in Table I. To guarantee the security and scalability of finalizer committees, we set the size of FCs the same as the shard size in the baseline. After that, we set the proposer shard size to the minimum value for better performance. We recommend that developers reduce FC and PS size for better performance without compromising security, according to Section 6.1. According to Table I, DL-Chain significantly reduces the PS size and increases the number of PS compared with the baseline. While PS sizes in DL-Chain are marginally larger than GearBox’s shard sizes, DL-Chain can accommodate more PSs. This is because the probability of system failure rises with the number of shard samples. In DL-Chain, the number of PSs is equivalent to the number of samples since we do not reshuffle corrupted shards. However, the number of shards in GearBox is lower than the number of samples, as GearBox requires multiple re-sampling attempts to obtain a single shard.
@@ 7.3 Parameter Settings
7.4 Transaction Throughput and Latency
We first compare the throughput (transaction per second, TPS) of DL-Chain and other systems at different network sizes, shown in Figures 5a and 5b. Compared with Harmony, DL-Chain has more PSs with smaller size, thus achieving more than 10 times the total TPS and 2 times the TPS for a single shard. Compared with GearBox, DL-Chain has more PSs and avoids network-wide consensus via scalable Dual-Layer Consensus, thus achieving up to 7 times the total TPS and 3 times the TPS for a single shard in a network size of 2,550.
We evaluate average transaction confirmation latency (i.e., the duration between a transaction starts to be processed until it is finalized, similar to previous works [19]), as shown in Figure 5c. Unlike the considerable gains in throughput, the transaction confirmation latency of DL-Chain is only a few seconds lower than that of Harmony. The reasons are that DL-Chain has a larger proportion of cross-shard transactions, and the confirmation of transactions awaits FCs’ finalization. However, DL-Chain reduces latency significantly compared with GearBox due to our Dual-Layer Consensus, which avoids network-wide consensus and ensures scalability and efficiency.
7.5 Breakdown of Block Latency
We analyze the block latency in DL-Chain’s three-phase Dual-Layer Consensus: block proposal in PSs, cross-layer verification, and block finalization in FCs. Unlike transaction confirmation latency, block latency represents the time from a proposer block’s proposal to its finalization. As depicted in Figure 5d, DL-Chain’s overall block latency is shorter than the baseline for three reasons. Firstly, fewer nodes participate in DL-Chain’s block proposal, reducing delay. Secondly, the system uses block headers for efficient cross-layer verification. Finally, while the FC sizes are comparable to the baseline shards, each FC’s consensus on metadata (proposer block hashes) rather than transactions accelerates block finalization.
@@ 7.5 Breakdown of Block Latency
7.6 Temporary Stagnation under Various Malicious Fractions
This experiment aims to assess DL-Chain’s ability to provide stable performance. We fix the network size as 640 and vary the fraction of malicious nodes to investigate its impact on system performance stability. This situation is common in real-world systems because adversaries can arbitrarily control the fraction of their malicious nodes to hinder system processes. Even in the absence of malicious nodes, the instability of the blockchain network environment can also lead to different percentages of nodes becoming unresponsive and hindering the system’s operation. We assume that the adversary silences 20% of the nodes at the beginning and silences another 5% of the nodes to initiate silent attacks to stagnate the system at the time $T=3.5$.
As shown in Figure 6a, the throughput of both systems drops at time $T$. For DL-Chain, the system wait longer to reach the quorum size for consensus due to additional silent nodes after time $T$, reducing efficiency. However, DL-Chain’s throughput is still more than 6 times the throughput of GearBox after time $T$. This is because most shards in GearBox lose liveness and cannot package any transactions. GearBox then underwent nearly 4 minutes of shard re-sampling and state migration involving 1 million Ethereum transactions [38]. However, the latency in real scenarios is even more significant than this and increases with state sizes, as shown in Figure 7. Finally, the throughput of GearBox is only partially recovered due to the increased shard size and reduced shard count.
As shown in Figure 6b, the latency of both systems increases at time $T$. For DL-Chain, the reason for increased latency is that it takes longer to reach a consensus within PSs or FCs when more nodes keep silent. However, the latency of GearBox significantly increases after time $T$. This is because all cross-shard transactions directed to the corrupted shards are stuck until the corrupted shard’s liveness is recovered.
@@ 7.6 Temporary Stagnation under Various Malicious Fractions
We evaluate the latency of state migration during reshuffling (only exists in GearBox) based on historical Ethereum transactions [38]. As shown in Figure 7, the more transactions executed, the longer time required for state migration. This is due to the greater involvement of accounts, resulting in a more intricate migration state. Note that the evaluation only involves 12 million Ethereum transactions for ten days [38], and the real-world situation of Ethereum, which has been running for several years, will involve more considerable delay. On the other hand, DL-Chain is immune to this time-consuming stagnation process.
8 Conclusion
This paper proposes DL-Chain, which effectively accommodates corrupted shards while preserving system security through its Dual-Layer Consensus. Our careful design strategies delegate the responsibility of liveness and transaction validity guarantee to each PS while tasking each FC with addressing the forking issue for PSs. These design approaches equipped DL-Chain with stable performance and reduced overhead. DL-Chain permits a larger proportion of malicious nodes in each PS (less than 1/2), enabling the secure configuration of smaller shards for enhanced concurrency. We demonstrate in our experiment that our system achieves stable high concurrency and a throughput improvement of 10 times.

### v2

@@ VII-A Implementation
We developed a StableShard prototype in Golang based on Harmony[Harmony], a prominent industrial permissionless blockchain sharding project that was once among the top 50 cryptocurrencies by market cap. The FBFT consensus protocol, proposed by Harmony, is used as our intra-shard consensus to ensure a fair comparison to demonstrate StableShard’s performance improvements. Harmony, a traditional blockchain sharding unable to tolerate corrupted shards, serves as our baseline protocol. We also implemented a GearBox (CCS 22) [gearbox] prototype as a state-of-the-art comparison. GearBox initially sets shards to a small size, prone to losing liveness. Upon detecting a corrupted shard, it reorganizes it until liveness is restored.
@@
VII-B Experimental Setup
We use a large-scale network of 2,550 nodes on 12 Amazon EC2 instances for testing. Each instance has a 128-core CPU and 50 Gbps network bandwidth, hosting up to 213 nodes. Docker containers and Linux Traffic Control managed inter-node communication, enforcing a 100 ms message delay and a 50 Mbps bandwidth limit per node. The experiment uses 512-byte transactions, and each block accommodates up to 4,096 transactions (i.e., 2MB blocks) as in existing work [rapidchain]. The transactions are based on historical data of Ethereum [xblock], in which the proportion of cross-shard transactions increases with the number of shards. The total fraction of malicious nodes is set at $f=1/4$, typical in practical network environments [buildingBlocks_survey].
VII-C Parameter Settings
In StableShard, proposer shard and finalizer committee sizes should be adjusted to guarantee a negligible system failure probability, as mentioned in Section VI-A. In the baseline (Harmony), we leverage the equations based on classical hypergeometric distribution in [rapidchain] to determine shard sizes. We leverage the function proposed by GearBox and conduct 10 million simulations from the beginning of the epoch to get the average number of shards for implementing GearBox. The negligible failure probability is less than $2^{-17}\approx 7.6\times 10^{-6}$, the same as many existing works [pyramid, li2022jenga], meaning that one failure occurs in about 359 years (with one-day epoch).
The parameter settings are shown in Table I. To guarantee the security and scalability of finalizer committees, we set the size of FCs the same as the shard size in the baseline. After that, we set the proposer shard size to the minimum value for better performance. We recommend that developers reduce FC and PS size for better performance without compromising security, according to Section VI-A. According to Table I, StableShard significantly reduces the PS size and increases the number of PS compared with the baseline. While PS sizes in StableShard are marginally larger than GearBox’s shard sizes, StableShard can accommodate more PSs. This is because the probability of system failure rises with the number of shard samples. In StableShard, the number of PSs is equivalent to the number of samples since we do not reshuffle corrupted shards. However, the number of shards in GearBox is lower than the number of samples, as GearBox requires multiple re-sampling attempts to obtain a single shard.
@@ VII-C Parameter Settings
VII-D Transaction Throughput and Latency
We first compare the throughput (transaction per second, TPS) of StableShard and other systems at different network sizes, shown in Figures 5a and 5b. Compared with Harmony, StableShard has more PSs with smaller size, thus achieving more than 10 times the total TPS and 2 times the TPS for a single shard. Compared with GearBox, StableShard has more PSs and avoids network-wide consensus via scalable DoLC protocol, thus achieving up to 7 times the total TPS and 3 times the TPS for a single shard in a network size of 2,550.
We evaluate average transaction confirmation latency (i.e., the duration between a transaction starts to be processed until it is finalized, similar to previous works [pyramid]), as shown in Figure 5c. Unlike the considerable gains in throughput, the transaction confirmation latency of StableShard is only a few seconds lower than that of Harmony. The reasons are that StableShard has a larger proportion of cross-shard transactions, and the confirmation of transactions awaits FCs’ finalization. However, StableShard reduces latency significantly compared with GearBox due to our DoLC protocol, which avoids network-wide consensus and ensures scalability and efficiency.
VII-E Breakdown of Block Latency
We analyze the block latency in StableShard’s three-phase DoLC protocol: block proposal in PSs, cross-layer verification, and block finalization in FCs. Unlike transaction confirmation latency, block latency represents the time from a proposer block’s proposal to its finalization. As depicted in Figure 5d, StableShard’s overall block latency is shorter than the baseline for three reasons. Firstly, fewer nodes participate in StableShard’s block proposal, reducing delay. Secondly, the system uses block headers for efficient cross-layer verification. Finally, while the FC sizes are comparable to the baseline shards, each FC’s consensus on metadata (proposer block hashes) rather than transactions accelerates block finalization.
@@ VII-E Breakdown of Block Latency
VII-F Temporary Stagnation under Various Malicious Fractions
This experiment aims to assess StableShard’s ability to provide stable performance. We fix the network size as 640 and vary the fraction of malicious nodes to investigate its impact on system performance stability. This situation is common in real-world systems because adversaries can arbitrarily control the fraction of their malicious nodes to hinder system processes. Even in the absence of malicious nodes, the instability of the blockchain network environment can also lead to different percentages of nodes becoming unresponsive and hindering the system’s operation. We assume that the adversary silences 20% of the nodes at the beginning and silences another 5% of the nodes to initiate silent attacks to stagnate the system at the time $T=3.5$.
As shown in Figure 6a, the throughput of both systems drops at time $T$. For StableShard, the system wait longer to reach the quorum size for consensus due to additional silent nodes after time $T$, reducing efficiency. However, StableShard’s throughput is still more than 6 times the throughput of GearBox after time $T$. This is because most shards in GearBox lose liveness and cannot package any transactions. GearBox then underwent nearly 4 minutes of shard re-sampling and state migration involving 1 million Ethereum transactions [xblock]. However, the latency in real scenarios is even more significant than this and increases with state sizes, as shown in Figure 7. Finally, the throughput of GearBox is only partially recovered due to the increased shard size and reduced shard count.
As shown in Figure 6b, the latency of both systems increases at time $T$. For StableShard, the reason for increased latency is that it takes longer to reach a consensus within PSs or FCs when more nodes keep silent. However, the latency of GearBox significantly increases after time $T$. This is because all cross-shard transactions directed to the corrupted shards are stuck until the corrupted shard’s liveness is recovered.
@@ VII-F Temporary Stagnation under Various Malicious Fractions
We evaluate the latency of state migration during reshuffling (only exists in GearBox) based on historical Ethereum transactions [xblock]. As shown in Figure 7, the more transactions executed, the longer time required for state migration. This is due to the greater involvement of accounts, resulting in a more intricate migration state. Note that the evaluation only involves 12 million Ethereum transactions for ten days [xblock], and the real-world situation of Ethereum, which has been running for several years, will involve more considerable delay. On the other hand, StableShard is immune to this time-consuming stagnation process.
VIII Conclusion
This paper presents StableShard, a sharded blockchain that achieves high concurrency with stable and scalable performance under corrupted shards. The key contribution is a PS-FC Division-of-Labor Collaboration (DoLC) protocol: many small proposer shards (PSs) maximize concurrency by proposing blocks quickly and ensuring transaction validity via quorum voting, while multiple large finalizer committees (FCs) provide scalable, header-only fork resolution and deterministic finality for multiple PSs without network-wide consensus. Crucially, FCs maintain stagnation-free, in-place liveness for faulty PSs through a cross-layer view-change mechanism and a jointly tuned quorum/committee configuration, thereby eliminating recovery-after-failure procedures that cause severe service interruption. By jointly parameterizing committee sizes and quorum thresholds, StableShard enables a Byzantine fraction approaching <1/2 in each PS while keeping each FC secure (<1/3) with overwhelmingly high probability, allowing many small PSs and multiple coexisting FCs for stable, scalable concurrency. A Harmony-based prototype and large-scale evaluation (up to 2,550 nodes) demonstrate up to 10x throughput improvement and significantly more stable concurrency under attacks compared with prior designs.
[Uncaptioned image] Mingzhe Li is currently a US-equivalent Assistant Professor with the School of Computing and Information Technology, Great Bay University. He received his Ph.D. degree from the Department of Computer Science and Engineering, Hong Kong University of Science and Technology in 2022. Prior to that, he received his B.E. degree from Southern University of Science and Technology. His research interests are mainly in blockchain sharding, consensus protocol, blockchain application, network economics, and crowdsourcing.
[Uncaptioned image] You Lin is currently a master candidate with Department of Computer Science and Engineering, Southern University of Science and Technology. He received his B.E. degree in computer science and technology from Southern University of Science and Technology in 2021. His research interests are mainly in blockchain, network economics, and consensus protocols.
[Uncaptioned image] Jin Zhang is currently an associate professor with Department of Computer Science and Engineering, Southern University of Science and Technology. She received her B.E. and M.E. degrees in electronic engineering from Tsinghua University in 2004 and 2006, respectively, and received her Ph.D. degree in computer science from Hong Kong University of Science and Technology in 2009. Her research interests are mainly in mobile healthcare and wearable computing, wireless communication and networks, network economics, cognitive radio networks and dynamic spectrum management.